OpenStudy (anonymous):

A car traveling 85 km/h strikes a tree. The front end of the car compresses and the driver comes to rest after traveling 0.80m. What is the average acceleration of the driver during the collision? Express the answer in terms of "g's", where 1.00g= 9.80m/s^2 The answer is -36 g's. I have no idea how I am not getting that. Any tips on setting this up?

OpenStudy (turingtest):

\[v_f^2=v_i^2+2ad\]in this case the acceleration is in the opposite direction as the initial velocity, and the final velocity is zero\[0=v_i^2-2ad\to a=\frac{v_i^2}{2d}\]divide this by g and you should get your answer.

OpenStudy (anonymous):

You have to care unit. car velocity unit km/h , g unit m/sn^2 and driver comes 0.80m